OpenStudy (anonymous):

If I saved $70 on an item that was 70% off. What was the original price?

OpenStudy (anonymous):

Can you try to write an equation? I will help by stating that your unknown x should be the original price

OpenStudy (anonymous):

I tried the percent deduction equation which is Sale Price=Original price - Original price(discount) but it wasn't correct.

OpenStudy (anonymous):

wat is ue name

zepdrix (zepdrix):

`saved $70` So the amount of the discount is $70. `the item was 70% off` The percent of discount, from the original price is 70%. We'll let the original price be \(\Large\bf\sf x\) as andras suggested. \[\Large\bf\sf $70\quad=\quad \text{%70}\cdot\;x\]70 dollars is the 70 percent discount of the original price. Then just solve for the original price.\[\Large\bf\sf 70\quad=\quad \frac{70}{100}x\]

zepdrix (zepdrix):

Hopefully I read that correctly :D lol These discount and pricing problems can be tricky.

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

just simple things, percent deduction equation sounds fancy :) 70%=70/100=0.7 So 70% off means 0.7 was taken from the full price 1-0.7=0.3 0.3x discounted price. saved 70 dollars means that the price I paid is x-70 0.3x=x-70

OpenStudy (anonymous):

idk that

OpenStudy (anonymous):

HAHA we do not agree :)

OpenStudy (anonymous):

about wat

zepdrix (zepdrix):

We have the same equation, yours is just fancier :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Hmmm it didn'y come out correct. What about if you save 100 on an item that was 30%

OpenStudy (anonymous):

didn't*

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

Well I think both of us did this wrong. 0.7x=x-70 0.3x=70 x=233.33333

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

70% of and 70% OFF is not the same thing! Confusing ..... marketing bull****

OpenStudy (whpalmer4):

Original price is $100. You paid 100%-70%=30% of $100, or $30, and thus saved $70.